# TabMGP: Martingale Posterior with TabPFN

### v1

@@ 5.1 Results
Table 2: Experiments on logistic regression: synthetic data with various link functions (top), real data with binary logistic regression (middle), and real data with multinomial logistic regression (bottom). We compare TabMGP (ours), BB, Copula, and Bayes. Reported values are coverage and the trace of the posterior covariance (median, in parentheses), computed over 100 repetitions. Target coverage is 0.95.
Marginal credible intervals.
In addition to the results of joint credible sets presented above, we also report the results of marginal credible intervals in Appendix H. We report both the coverage and the Winkler score (Winkler, 1972), a widely used scoring rule for assessing univariate interval estimates. By this metric, TabMGP tends to perform better for logistic regression, whereas BB performs better for linear regression despite being severely under-covered most of the time. Across the 293 regression coefficients considered, TabMGP is preferred 155 times, BB 125 times, Copula 11 times, and Bayes 2 times.
Computation time.
@@
The main computational cost of martingale posteriors is forward sampling, which is readily parallelisable. Per sample, BB is fastest (near-instantaneous); Copula takes 10–20s; and TabMGP takes about 70s (binary/categorical response) or 200s (continuous response). The NUTS algorithm (Bayes) runs in under 60s. BB, Copula, and Bayes are implemented in compiled JAX, whereas TabMGP currently runs in uncompiled PyTorch. We expect this gap to narrow with further code optimisation. All experiments were conducted on an Nvidia L40s.
@@ 6 Conclusion
We introduced TabMGP, a new construction of the martingale posterior that uses a pre-trained tabular foundation model, TabPFN, as its predictive rule. This approach circumvents the usual challenges of prior specification in classical Bayesian inference and the need to design bespoke predictive rules in existing martingale posterior frameworks. By utilising the in-context learning capabilities of transformers, TabMGP provides an off-the-shelf tool for statistical inference. Across a wide range of synthetic and real-world datasets, our empirical evaluation shows that TabMGP delivers reliable uncertainty quantification.
Our findings also expose limitations in the current theoretical understanding of martingale posteriors. Existing analyses emphasise sufficient conditions such as the martingale or a.c.i.d. properties, yet TabPFN fails both while still yielding posteriors that converge and achieve near-nominal coverage in practice. This suggests that the present theory is overly restrictive and does not yet capture the range of predictive rules that work well empirically. Developing weaker and more realistic conditions is an important direction for future research, particularly as martingale posteriors are likely to be combined with an increasing variety of black-box predictive models. Strengthening the theoretical foundations of this marriage will be crucial for ensuring that such methods can be used with confidence in practice.
Beyond its role in the post-Bayes expansion, our perspective is also potentially relevant for deep learning interpretability. A pretrained transformer, when used through in-context learning, implicitly defines a predictive rule. The Bayesian predictive framework raises the intriguing possibility that if this rule induces asymptotic exchangeability, then de Finetti’s representation theorem applies and the transformer can be viewed as carrying a hidden prior–likelihood pair (Fortini and Petrone, 2025). Thus the Bayesian predictive framework could provide a new lens on the internal structure of foundation models.
@@ Appendix A Review of TabPFN
A precursor to TabPFN can be found in Müller et al. (2022), where the transformer is simply called a prior-data-fitted network (PFN). The premise is predicated that shows the Bayes posterior predictive density is optimal in an average sense. We present this in the unsupervised setting to keep the notation lightweight on a result [^1] Though the work Müller et al. (2022) does not appear to be aware of this result in Aitchison (1975).
@@
Aitchison (1975) then shows that the predictive density $q$ that minimizes this average KL divergence in (6) is none other than the Bayesian posterior predictive density (PPD) corresponding to the likelihood-prior pair in (5):
@@
Later developments led to the TabPFN Hollmann et al. (2022; 2025) which trains on a vast amount of synthetic data generated from a very large array of likelihood-prior pairs, rather than a single pair as in the original PFN work. This gives TabPFN the qualifier foundational.
@@
We evaluate this condition on a dataset generated as $y_{i}\overset{i.i.d.}{\sim}\operatorname{Bernoulli}(\operatorname{logistic}(x^{\top}_{i}\beta_{0}))$, for $i=1,\dots,100$, with some fixed $\beta_{0}\in\mathbb{R}^{2}$. Our empirical results (Figure 3) indicate that the predictive rule induced by TabPFN does not strictly satisfy the a.c.i.d. condition. This negative result, however, does not imply non-convergence to $F_{\infty}$ for TabMGP, as the a.c.i.d. condition is only sufficient, not necessary.
Figure 3: Cumulative sum (over NN) of the $L_{1}$-distance between $\mathbb{E}_{y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]$ and $p_{i}(y\mid x^{\star})$ across eight choices of $x^{\star}$. Ideally, this cumulative sum converges to satisfy the a.c.i.d. condition.
@@
For completeness, we further demonstrate that TabMGP can concentrate close to $\theta_{0}$ as the sample size $n$ grows. To illustrate this, we use a regression dataset, since the sum of the sample size needed for meaningful concentration and the forward sampling steps required for convergence in classification exceeds TabPFN’s 10,000-context length limit. The regression dataset is defined as $y_{i}\overset{i.i.d.}{\sim}{\mathcal{N}}(x^{\top}_{i}\beta_{0},0.1^{2}),\,i=1,\dots,n$, for some fixed $\beta_{0}\in\mathbb{R}^{2}$, with the loss function $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$. Ideally, TabMGP should concentrate on $\theta_{0}=[0,\beta_{0}^{\top}]$ as $n$ increases. Figure 4 confirms that TabMGP indeed concentrates around $\theta_{0}$.
Figure 4: Concentration of TabMGP. The red vertical line indicates the population risk minimiser $(0,\beta_{0}^{\top})$.
@@ E.1 Posterior constructions
Settings of Bayes.
We approximate the posterior using a No-U-turn sampler (NUTS, Hoffman and Gelman, 2014) with 1000 warm-up steps, followed by 2000 sampling steps across 4 chains with different initialisations, for a total of 8000 samples.
Settings for all martingale posteriors.
We run 2000 forward-sampling steps for both BB and Copula. For TabMGP, we find that $N=n+500$ is sufficient for $\theta(F_{N})$ to converge (Figure 1 and Appendix F) and use only 500 forward-sampling steps. We approximate all martingale posterior constructions with 100 samples.
Settings of TabMGP.
The logits of the predictive rules are obtained from TabPFNClassifier.predict_proba for logistic regression and TabPFNRegressor.predict(output_type = "full") for linear regression. By default, these logits are scale by a temperature of 0.9. For our experiments, we use the untempered logits as our predictive rules, i.e., setting the temperature to 1. By construction, TabPFN is not permutation invariant to the order of the dimensions of labels $y$ and features $x$. The default implementation of TabPFN uses an ensemble of predictions from randomly permuting the dimensions of labels and features. The resulting probability mass functions are then averaged to get the final predictive mass function. We use the default number of ensemble in version 2.0.6, which is 4 for classifiers and 8 for regressors.
@@
Settings of Copula.
We use the copula construction in Fong et al. (2023, §4.4.2 and §4.4.3) with the bandwidth fixed to 0.8, since for our experiments the marginal-likelihood-based tuning procedure of Fong et al. (2023) produces values that are too small and delivers poor results. Instead, we estimate
@@ E.2 Specification of the loss function
We identify the collinear features as those with the largest loading in the least important principal component of the population data, removing them one at a time until the computation of $\theta(F_{0})$ is numerically stable. The resulting parameter dimension, $p\coloneq\dim(\theta)$, and the features used in the functional are listed in Table 3 and Table 4, respectively.
@@ E.3 Data setups
Synthetic data.
The 11 synthetic setups are adapted from Wu and Martin (2023, Section 5): the features $x_{i}$ are i.i.d. and uniformly sampled from the hypercube $1,1]^{10}$. The true coefficients $\beta_{0}$ are sampled once from $2,3]^{10}$ and held fixed across the 100 repetition of $z_{1:n}$. The response $y\mid x$ is generated as follows:
  • Continuous i.i.d. response (4 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta_{0}+\epsilon_{i},\,i=1,\ldots,n$. The errors $\epsilon_{i}$ are either i.i.d. ${\mathcal{N}}(0,1)$ or Student-$t$ with 5, 4, or 3 degrees of freedom. Note that $\theta(F_{0})$ coincides with $(0,\beta_{0}^{\top})$ in the case of i.i.d. $N(0,1)$ error.
  • Continuous response with heteroscedastic errors (3 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta_{0}+\epsilon_{i},\,i=1,\ldots,n$, where $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma_{i}^{2})$. The standard deviation of the error $\sigma_{i}$ is determined by the first coordinate of $x_{i}$:
@@
  • Binary response (4 setups): We generate $n=100$ samples from $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta_{0})),\,i=1,\ldots,n$, with different link functions $L$. The link is either the logistic function:
@@
    where $\Phi(\cdot\mid m,1)$ denotes the distribution function of a Gaussian with mean $m$ and unit variance, and $a\in\{0,-1,-2\}$. Note that $\theta(F_{0})$ coincides with $(0,\beta_{0}^{\top})$ in the case of logistic link.
Real data.
We apply train_test_split in scikit-learn (Pedregosa et al., 2011) on the whole dataset, and use the training split as the dataset $z_{1:n}$. We repeat this procedure 100 times to generate repetitions of $z_{1:n}$. For datasets with highly skewed responses or features, we apply stratified sampling on these variables. For yeast and wine, which have highly skewed multi-class categorical responses, we also remove classes with fewer than 3% of observations. Multinomial logistic regression on a small dataset with an extremely skewed response remains difficult with our method, as the design matrix tends to be ill-conditioned when applying the Bayesian bootstrap on $x$.
@@
$z_{1:n}$ Data ID $n$ $p$ $n/p$
Population
size
Num.
classes
Num. cont.
features
Num. cat.
features
@@ E.4 Constructions of credible sets
Joint credible set
Let $\widehat{\mu}$ and $\widehat{\Sigma}$ denote the empirical mean and covariance estimated from the posterior draws $\{\theta^{(l)}\}_{l=1}^{L}$. For martingale posteriors, we use a diagonal covariance structure $\widehat{\Sigma}=\operatorname{diag}(\widehat{\sigma}^{2}_{1},\ldots,\widehat{\sigma}^{2}_{p})$ due to the limited number of posterior samples available for estimating a full covariance matrix. Our credible set is then:
@@
Marginal credible interval.
Let $\theta_{j}$ be the $j$-th dimension of $\theta$. The $(1-\alpha)$ credible interval is computed from the empirical quantiles of its posterior draws:

### v2

@@ 5.1 Results
Table 2: Experiments on logistic regression: synthetic data with various link functions (top), real data with binary logistic regression (middle), and real data with multinomial logistic regression (bottom). The rows corresponding to real data are sorted by $n/p$ ratio, i.e, bottom rows are harder. We compare TabMGP (ours), BB, Copula, Bayes and Asymptotic. Reported values are coverage (Rate) and the median of trace of the posterior covariance (Size), computed over 100 repetitions. Target coverage is 0.95.
TabMGP $B$ Copula Bayes Asymptotic
Setup Rate Size Rate Size Rate Size Rate Size Rate Size
Logistic 0.84 1.30 0.83 1.96 1.00 2.72 0.53 0.78 0.96 1.53
GMM(0) 0.91 2.44 0.91 4.14 1.00 4.01 0.60 1.34 0.98 2.63
GMM(-1) 0.83 1.27 0.86 1.95 1.00 2.64 0.58 0.79 0.99 1.54
GMM(-2) 0.78 0.88 0.78 1.36 1.00 2.08 0.52 0.57 1.00 1.13
rice 1.00 3.74 0.88 7.35 1.00 28.41 0.91 2.87 1.00 5.92
sepsis 0.95 0.69 1.00 2.42 1.00 10.00 0.98 1.18 1.00 2.42
banknote 0.99 1.16 1.00 1.48 1.00 1.53 0.99 0.64 1.00 1.28
mozilla 0.96 1.42 0.97 2.04 1.00 2.68 0.87 0.65 0.93 1.31
skin 1.00 1.39 1.00 1.84 1.00 1.80 1.00 0.57 1.00 1.16
blood 0.87 1.07 0.95 2.51 1.00 5.40 0.84 0.86 0.99 1.71
phoneme 1.00 1.32 0.95 1.92 1.00 5.62 0.73 0.83 1.00 1.66
telescope 0.99 5.96 0.99 10.66 1.00 50.69 0.67 3.50 0.96 7.10
yeast 0.97 11.10 0.68 25.13 0.33 10.82 0.88 19.22
wine 0.93 11.04 0.79 39.75 0.09 12.51 0.91 26.25
Marginal credible intervals. In addition to the results of joint credible sets presented above, we also report the results of marginal credible intervals in Appendix H. We report both the coverage and the median width of the credible intervals.
@@
Computation time. The main computational cost of martingale posteriors is forward sampling, which is readily parallelisable. Per sample, BB and Asymptotic are the fastest (near-instantaneous); Copula takes 10–20s; and TabMGP takes about 70s (binary/categorical response) or 200s (continuous response). However, despite being the most computationally expensive method, TabMGP requires minimal manual intervention compared to the others. The NUTS algorithm (Bayes) runs in under 60s. BB, Copula, Bayes, and Asymptotic are implemented in compiled JAX, whereas TabMGP currently runs in uncompiled PyTorch. We expect this gap to narrow with further code optimisation. All experiments were conducted on an Nvidia L40s.
@@ 6 Conclusion
While bespoke Bayesian modelling remains the gold standard when a well-specified likelihood–prior pair is available, tabular foundation models like TabPFN are increasingly used in practice. In scientific workflows, prediction is usually a means to an end: one ultimately wants uncertainty about a scientific estimand $\theta$. However, TabPFN does not by itself return posterior uncertainty for $\theta$. In TabPFN, the carrier of epistemic uncertainty is the predictive rule itself and the associated random limiting measure $F_{\infty}$, so the natural inference target is the functional $\theta(F_{\infty})$. This motivates a framework that returns posterior uncertainty for $\theta(F_{\infty})$ using only queries to TabPFN’s predictive outputs.
For this purpose, we introduced TabMGP. The procedure requires neither likelihood–prior specification, as in classical Bayesian inference, nor bespoke predictive rules or hyperparameter tuning, as in existing martingale posterior frameworks, relying instead solely on the predictive rule induced by TabPFN. We evaluated TabMGP across a large collection of synthetic setups and real-world datasets, and compared extensively to existing martingale posterior constructions as well as standard Bayesian baselines. Across these experiments, TabMGP was consistently favourable in terms of frequentist behaviour, delivering stable posteriors with near-nominal coverage. Our empirical evaluation shows that TabMGP provides reliable uncertainty quantification.
Our findings also expose limitations in the current theoretical understanding of martingale posteriors. Existing analyses emphasise sufficient conditions such as the martingale or a.c.i.d. properties; in our experiments, finite-horizon diagnostics do not certify these sufficient conditions, yet the resulting posteriors maintain empirical stability and near-nominal coverage. This suggests that the present theory may be overly restrictive and does not yet capture the range of predictive rules that work well empirically. Developing weaker and more realistic conditions is an important direction for future research, particularly as martingale posteriors are likely to be combined with an increasing variety of black-box predictive models. More broadly, deep learning theory continues to lag far behind practice: for predictors like TabPFN, we should not expect to be able to theoretically verify that any proposed condition holds, so progress will also require a battery of theory-guided empirical diagnostics to test these conditions.
@@ Appendix A Review of TabPFN
A precursor to TabPFN appears in Müller et al. (2022), where the transformer is referred to as a prior-data-fitted network (PFN). The underlying premise is that the Bayesian posterior predictive distribution (PPD) is optimal in an average sense. We present this result in the unsupervised setting to keep the notation lightweight, drawing on a result from Aitchison (1975) [^1] The work of Müller et al. (2022) does not appear to be aware of this result..
@@
Aitchison (1975) then shows that the predictive density $q$ that minimises this average KL divergence in (6) is none other than the Bayesian PPD corresponding to the likelihood–prior pair in (5):
@@
Later developments led to the TabPFN (Hollmann et al., 2022; 2025) which trains on a vast amount of synthetic data generated from a very large array of likelihood–prior pairs (e.g., structural causal models and Bayesian neural network models), rather than a single pair as in the original PFN work. This gives TabPFN the qualifier foundational.
@@
We evaluate this condition on an initial dataset of size $n=100$ generated as follows: $x_{i}$ are drawn independently and uniformly from the two-dimensional cube $1,1]^{2}$, and $y_{i}\overset{\mathrm{i.i.d.}}{\sim}\operatorname{Bernoulli}(\operatorname{logistic}(x_{i}^{\top}\beta^{\star}))$, for $i=1,\ldots,n$, with a fixed $\beta^{\star}\in\mathbb{R}^{2}$. We consider 10 different values of $x^{\star}$, randomly selected from the initial sequence $x_{1},\ldots,x_{n}$. Due to computational constraints, we evaluate (11) only on a sparse grid $i\in\{100,120,140,\ldots,1100\}$ and use the trapezoidal rule to approximate the partial sum.
Our empirical results (Figure 5) suggest that the predictive rule induced by TabPFN exhibits a lack of strict adherence to the a.c.i.d. condition as $N$ increases. However, at these scales ($N>1000$), any measurable drift falls below the numerical resolution of the model, making it impossible to distinguish genuine theoretical departures from the cumulative numerical instability inherent to long-horizon autoregressive inference. This empirical ambiguity, however, does not imply non-convergence to $F_{\infty}$ for TabMGP, as the a.c.i.d. condition is sufficient but not necessary. Moreover, the observed divergence may be attributable to the finite numerical precision of TabPFN. In particular, both $\mathbb{E}_{Y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]$ and $p_{i}(y\mid x^{\star})$ are expected to be accurate only up to machine precision, so that their difference behaves as random numerical noise for sufficiently large $N$, and the partial sum is infeasible to compute in this regime.
Figure 5: Cumulative sum (over NN) of the $L_{1}$-distance between $\mathbb{E}_{y_{i+1}}[p_{i+1}(y\mid x^{\star})\mid z_{1:i}]$ and $p_{i}(y\mid x^{\star})$ across ten choices of $x^{\star}$. Ideally, this cumulative sum converges to satisfy the a.c.i.d. condition.
@@
For completeness, we further demonstrate that TabMGP can concentrate close to $\theta(F^{\star})$ as the sample size $n$ grows. To illustrate this, we use a regression dataset, since the sum of the sample size needed for meaningful concentration and the forward sampling steps required for convergence in classification exceeds TabPFN’s 10,000-context length limit. The regression dataset is defined as $y_{i}\overset{i.i.d.}{\sim}{\mathcal{N}}(x^{\top}_{i}\beta^{\star},0.1^{2}),\,i=1,\dots,n$, for some fixed $\beta^{\star}\in\mathbb{R}^{2}$, with the loss function $\ell(x,y,\theta)=(y-[1\ x^{\top}]\theta)^{2}$. Ideally, TabMGP should concentrate on $\theta(F^{\star})=[0,{\beta^{\star}}^{\top}]$ as $n$ increases. Figure 2 confirms that TabMGP indeed concentrates around $\theta(F^{\star})$.
@@ E.1 Posterior constructions
Asymptotic approximation of the Bayesian posterior. Let $\widehat{\theta}_{n}=\operatorname*{arg\,min}_{\theta}\sum^{n}_{i=1}\ell(z_{i},\theta)$ and $\widehat{H}_{n}=\nabla^{2}_{\theta}\sum^{n}_{i=1}\ell(z_{i},\theta)|_{\theta=\widehat{\theta}_{n}}$. We use ${\mathcal{N}}(\widehat{\theta}_{n},\widehat{H}_{n}^{-1})$ as the asymptotic approximation of the posterior for both constructing the credible region and using it as the prior for the classical Bayes baseline.
Settings of Bayes. We approximate the posterior using a No-U-turn sampler (NUTS, Hoffman and Gelman, 2014) with 1000 warm-up steps, followed by 2000 sampling steps across 4 chains with different initialisations, for a total of 8000 samples.
Settings for all martingale posteriors. We run 2000 forward-sampling steps for both BB and Copula. For TabMGP, we find that $N=n+500$ is sufficient for $\theta(F_{N})$ to converge (Figure 3 and Appendix F) and use only 500 forward-sampling steps. We approximate all martingale posterior constructions with 100 samples.
Settings of TabMGP. The logits of the predictive rules are obtained from TabPFNClassifier.predict_proba for logistic regression and TabPFNRegressor.predict(output_type = "full") for linear regression. By default, these logits are scale by a temperature of 0.9. For our experiments, we use the untempered logits as our predictive rules, i.e., setting the temperature to 1. By construction, TabPFN is not permutation invariant to the order of the dimensions of labels $y$ and features $x$. The default implementation of TabPFN uses an ensemble of predictions from randomly permuting the dimensions of labels and features. The resulting probability mass functions are then averaged to get the final predictive mass function. We use the default number of ensemble in version 2.0.6, which is 4 for classifiers and 8 for regressors.
@@
Settings of Copula. We use the copula construction in Fong et al. (2023, §4.4.2 and §4.4.3) with the bandwidth fixed to 0.8, since for our experiments the marginal-likelihood-based tuning procedure of Fong et al. (2023) produces values that are too small and delivers poor results. Instead, we estimate
@@ E.2 Specification of the loss function
We identify the collinear features as those with the largest loading in the least important principal component of the population data, removing them one at a time until the computation of $\theta(F^{\star})$ is numerically stable. The resulting parameter dimension, $p\coloneq\dim(\theta)$, and the features used in the functional are listed in Table 3 and Table 4, respectively.
@@ E.3 Data setups
Synthetic data. The 11 synthetic setups are adapted from Wu and Martin (2023, Section 5): the features $x_{i}$ are i.i.d. and uniformly sampled from the hypercube $1,1]^{10}$. The true coefficients $\beta^{\star}$ are sampled once from $2,3]^{10}$ and held fixed across the 100 repetition of $z_{1:n}$. The response $y\mid x$ is generated as follows:
  • Continuous i.i.d. response (4 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i},\,i=1,\ldots,n$. The errors $\epsilon_{i}$ are either i.i.d. ${\mathcal{N}}(0,1)$ or Student-$t$ with 5, 4, or 3 degrees of freedom. Note that $\theta(F^{\star})$ coincides with $(0,{\beta^{\star}}^{\top})$ in the case of i.i.d. $N(0,1)$ error.
  • Continuous response with heteroscedastic errors (3 setups): We generate $n=20$ samples from $y_{i}=x^{\top}_{i}\beta^{\star}+\epsilon_{i},\,i=1,\ldots,n$, where $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma_{i}^{2})$. The standard deviation of the error $\sigma_{i}$ is determined by the first coordinate of $x_{i}$:
@@
  • Binary response (4 setups): We generate $n=100$ samples from $y_{i}\mid x_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Bernoulli}(L(x^{\top}_{i}\beta^{\star})),\,i=1,\ldots,n$, with different link functions $L$. The link is either the logistic function:
@@
    where $\Phi(\cdot\mid m,1)$ denotes the distribution function of a Gaussian with mean $m$ and unit variance, and $a\in\{0,-1,-2\}$. Note that $\theta(F^{\star})$ coincides with $(0,{\beta^{\star}}^{\top})$ in the case of logistic link.
Real data. We apply train_test_split in scikit-learn (Pedregosa et al., 2011) on the whole dataset, and use the training split as the dataset $z_{1:n}$. We repeat this procedure 100 times to generate repetitions of $z_{1:n}$. For datasets with highly skewed responses or features, we apply stratified sampling on these variables. For yeast and wine, which have highly skewed multi-class categorical responses, we also remove classes with fewer than 3% of observations. Multinomial logistic regression on a small dataset with an extremely skewed response remains difficult with our method, as the design matrix tends to be ill-conditioned when applying the Bayesian bootstrap on $x$.
@@
Table 3: Details of the real-world datasets.
$z_{1:n}$ Data ID $n$ $p$ $n/p$ Population size Num. classes Num. cont. features Num. cat. features
@@ E.4 Constructions of credible sets
Joint credible set Let $\widehat{\mu}$ and $\widehat{\Sigma}$ denote the empirical mean and covariance estimated from the posterior draws $\{\theta^{(l)}\}_{l=1}^{L}$. For martingale posteriors, we use a diagonal covariance structure $\widehat{\Sigma}=\operatorname{diag}(\widehat{\sigma}^{2}_{1},\ldots,\widehat{\sigma}^{2}_{p})$ due to the limited number of posterior samples available for estimating a full covariance matrix. Our credible set is then:
@@
Marginal credible interval. Let $\theta_{j}$ be the $j$-th dimension of $\theta$. The $(1-\alpha)$ credible interval is computed from the empirical quantiles of its posterior draws: